My problem is C squared equals 8x squared(substituted for A) plus 15x squared(once again substituted)
Since the question is an odd one, i got my answer of 23x squared but the book got something different...
The book got 17x
i do not understand how they got that.

Answer by oscargut(2103)

C^2=(8x)^2+(15x)^2
then
C^2= 64x^2+225x^2
C^2=289x^2
then
C= sqrt(289)x=17x